Solve the inequality.
(c+4)(c-2)(c-7)>0

Answer by RAY100(1637)

(c+4)(c-2)(c-7) > 0
.
lets solve as equality first
.
(c+4)(c-2)(c-7) =0
.
c+4=0,,,,c=-4
.
c-2=0,,,,c=2
.
c-7=0,,,,c=7
.
in order,,,,,-4,,,,2,,,,,,7
.
now lets check numbers before and after each of these against original inequality
.
for x less than (-4),,, let x=-5,,, -84 >0,,,, not true
.
for -4 less than x less than 2,,,let x=0,,,,,56>0,,,,,ok
.
for 2 less than x less than 7,,,,,let x= 3,,,,,-28>0,,,, not true
.
for x>7,,,,,let x = 8,,,,,72>0,,,ok
.
therefore ,,,,-4less than x less than 2,,,,x>7,,,are answers